U.S. Steel is considering a plant expansion to produce austenitic, precipitation hardened, duplex, and martensitic stainless steel round bars that is expected to cost $14 million now and another $10 million 1 year from now. If total operating costs will be $1.4 million per year starting 1 year from now, and the estimated salvage value of the plant is virtually zero, how much must the company make annually in years 1 through 9 to recover its investment plus a return of 24% per year? The company must make $ million annually in years 1 through 9 to recover its investment plus a return of 24% per year.
Answer #1

Present worth (PW) of costs ($ million) = 14 + 10 x P/F(24%, 1) + 1.4 x P/A(24%, 9)

= 14 + 10 x 0.8065** + 1.4 x 3.5655** = 14 + 8.06 + 4.9917 = 27.06

Annual income ($ million) = PW / P/A(24%, 9) = 27.06 / 3.5655** = 7.59

**From P/F and P/A Factor tables